OpenStudy (anonymous):

Expand and evaluate the power. 28 A. 16 B. 256 C. 512 D. 1024

OpenStudy (anonymous):

I think it's B or C. I need to verify.

OpenStudy (freethinker):

28 what?

OpenStudy (anonymous):

2^8 sorry

OpenStudy (freethinker):

oh

OpenStudy (anonymous):

the eight power of 2 lol

OpenStudy (freethinker):

can't you multiply by yourself? 2^8 = 2*2*2*2*2*2*2*2

OpenStudy (anonymous):

yah lol

OpenStudy (anonymous):

8x8= 64

OpenStudy (freethinker):

2*2 = 4 4*2 = 16 16*2 = 32 32*2 = 64 64*2 = 128

OpenStudy (anonymous):

so the answer is B

OpenStudy (anonymous):

thnx

OpenStudy (freethinker):

learn multiplication, kid

OpenStudy (anonymous):

2^ 8.... It means 2 times itself 8 times. 2*2*2*2*2*2*2*2